Exemplar Solution

Get insights from 1.3k questions on Exemplar Solution, answered by students, alumni, and experts. You may also ask and answer any question you like about Exemplar Solution

  • 1.3k

    Questions

  • 0

    Discussions

  • 0

    Active Users

  • 0

    Followers

Have a question about Exemplar Solution? Ask our experts

Ask Your Question

Keep it short & simple. Type complete word. Avoid abusive language.

New answer posted•a year ago
S
a year ago

This is a Fill in the blanks Type Questions as classified in NCERT Exemplar

L e t     E     b e     t h e     e v e n t     t h a t     t h e     s t u d e n t     w i l l     p a s s a n d     F     b e     t h e     e v e n t     t h a t     h e     w i l l     g e t     c o m p a r t m e n t ∴ P ( E ) = 0 . 7 3 , P ( F ) = 0 . 1 3     a n d     P ( E ∪ F ) = 0 . 9 6 ∴                   P ( E ∪ F ) = P ( E ) + P ( F ) − P ( E ∩ F )                                                                     = 0 . 7 3 + 0 . 1 3 − 0                       [ ? P ( E ∩ F ) = 0 ]                                                                     = 0 . 8 6 B u t         P ( E ∪ F ) = 0 . 9 6 H e n c e ,     t h e     g i v e n     s t a t e m e n t     i s     ' F a l s e ' .

Thumbs Up IconUpvote Thumbs Down Icon
New answer posted•a year ago
S
a year ago

This is a Fill in the blanks Type Questions as classified in NCERT Exemplar

G i v e n     t h a t :                                                                   P ( t o     s e e     g i r a f f e e ) = 0 . 7 2                                                                               P ( t o     s e e     b e a r ) = 0 . 8 4 P ( t o     s e e     b o t h     g i r a f f e e     a n d     b e a r ) = 0 . 5 2 ∴                         P ( t o     s e e     g i r a f f e e     o r     b e a r ) = P ( t o     s e e     g i r a f f e e ) + P ( t o     s e e     b e a r ) − P ( t o     s e e     b o t h )                                                                                                                                                   = 0 . 7 2 + 0 . 8 4 − 0 . 5 2                                                                                                                                                   = 1 . 0 4     w h i c h     i s     n o t     p o s s i b l e . H e n c e ,     t h e     g i v e n     s t a t e m e n t     i s     ' F a l s e ' .

Thumbs Up IconUpvote Thumbs Down Icon
New answer posted•a year ago
S
a year ago

This is an Objective Type Questions as classified in NCERT Exemplar

If  M  and  N  are  any  two  events,   then          P (M∪N)=P (M)+P (N)−P (M∩N)Hence,   the  correct  option  is   (b).

Thumbs Up IconUpvote Thumbs Down Icon
New answer posted•a year ago
S
a year ago

This is an Objective Type Questions as classified in NCERT Exemplar

Given  that:           P(A∪B)=0.6                                    P(A∩B)=0.2∴          P(A∪B)=P(A)+P(B)−P(A∩B)⇒                       0.6=P(A)+P(B)−0.2⇒P(A)+P(B)=0.6+0.2=0.8and        1−P(A¯)+1−P(B¯)=0.8⇒     P(A¯)+P(B¯)=2−0.8=1.2Hence,  the  correct  option  is  (c).

Thumbs Up IconUpvote Thumbs Down Icon
New answer posted•a year ago
S
a year ago

This is an Objective Type Questions as classified in NCERT Exemplar

Given  that:           P(A  fails)=0.2                                    P(B  fails)=0.3∴        P(either  A  or  B  fails)≤P(A  fails)+P(B  fails)                                                              ≤0.2+0.3                                                              ≤0.5Hence,  the  correct  option  is  (c).

Thumbs Up IconUpvote Thumbs Down Icon
New answer posted•a year ago
S
a year ago

This is an Objective Type Questions as classified in NCERT Exemplar

Total  number  of  alphabets  in  probability=11Number  of  vowels=4∴   Required probability=411Hence,   the  correct  option  is   (b).

Thumbs Up IconUpvote Thumbs Down Icon
New answer posted•a year ago
S
a year ago

This is an Objective Type Questions as classified in NCERT Exemplar

If  all  the  girls  sit  together,  then  we  consider  it  as  1  group∴  Total  number  of  arrangement  of  6+1=7  persons  in  a  row=7!and  the  girls  also    their  places  with  6!  ways.∴    Required probability=6! 7!12!=6×5×4×3×2×7!12×11×10×9×8×7!=1132Hence,  the  correct  option  is  (c).

.

Thumbs Up IconUpvote Thumbs Down Icon
New answer posted•a year ago
S
a year ago

This is an Objective Type Questions as classified in NCERT Exemplar

G i v e n     t h a t :                 P ( A ∪ B ) = P ( A ∩ B ) ⇒                                                                         P ( A ) + P ( B ) − P ( A ∩ B ) = P ( A ∩ B ) ⇒ [ P ( A ) − P ( A ∩ B ) ] + [ P ( B ) − P ( A ∩ B ) ] = 0 B u t     P ( A ) − P ( A ∩ B ) ≥ 0                   [ ? P ( A ∩ B ) ≤ P ( A )     o r     P ( B ) ]                                                     … ( i ) a n d     P ( B ) − P ( A ∩ B ) ≥ 0                                                                                                                                                                                                           … ( i i ) F r o m     e q n . ( i )     a n d     ( i i )     w e     g e t                           P ( A ) = P ( B ) H e n c e ,     t h e     c o r r e c t     o p t i o n     i s     ( a ) .

Thumbs Up IconUpvote Thumbs Down Icon
New answer posted•a year ago
S
a year ago

This is an Objective Type Questions as classified in NCERT Exemplar

F o r     m u t u a l l y     e x c l u s i v e     e v e n t s ,                     P ( A ∩ B ) = 0 ∴               P ( A ∪ B ) = P ( A ) + P ( B )                 [ ? P ( A ∩ B ) = 0 ] ⇒                 P ( A ) + P ( B ) ≤ 1 ⇒ P ( A ) + 1 − P ( B ¯ ) ≤ 1                                                 [ ? P ( B ) = 1 − P ( B ¯ ) ] ⇒                 P ( A ) − P ( B ¯ ) ≤ 0 ⇒                                                 P ( A ) ≤ P ( B ¯ ) H e n c e ,     t h e     c o r r e c t     o p t i o n     i s     ( a ) .

Thumbs Up IconUpvote Thumbs Down Icon